Monday, July 23, 2012

Day 57, 7/23/2012 -- BEAMS

Last night was great success for overnight team. Make many beams, is good quality!

Ahem...yeah, so apparently last night went pretty well! This morning I came in and climbed the water tower to bring the source down while Franky did some noise tests. I ended up accidentally unplugging some of the lighthouse lights, so I had to climb the water tower again in the afternoon to fix that (loose cables). Franky warmed the focal plane to recover the wafers for one of the crates that overheated yesterday, and it is now cooling so as to be ready for more scans tonight. The agenda tonight is to the first attempt at our absolute polarization rotation measurement.

In comings and goings, this morning Amber and Britt headed back to New York while Michele is being picked up by KyleH at the airport. And, sadly (for you!), I'm taking a break from Palestine for a little while, so the updates will pause here until I return.

I don't have any EBEX pictures today, but I took a couple more BLAST and SuperTIGER pictures, which you can see in their respective albums:
https://picasaweb.google.com/104253244018605213307/BLAST2012
https://picasaweb.google.com/104253244018605213307/SuperTIGER2012

Sunday, July 22, 2012

Day 56, 7/22/2012 -- Sunday sun day

Various things done today -- Britt and I got the two solar arrays fully populated with solar panels, we cycled the fridges, we got the bolo system connected to its flight power system with a battery (built-in UPS!), filled helium, and the night shift is getting ready to take more water tower scans overnight tonight.

In the morning, I had to climb the water tower to retrieve a piece of equipment that I forgot to bring down yesterday; it had to weather the storm (covered by a plastic box) and I had to make sure it still worked. And, despite my bumbling, it did, even after I accidentally connected the +12V power to the output and tried to measure the output signal from the 12V input power connector. Oops.

The overnight team had a dramatic victory last night, though. As I mentioned in yesterday's post, there were thunderstorms in Palestine overnight. The power at the high bay often goes out during these thunderstorms, and if the bolometer system loses power, we lose all of our bolometer biases, the bolometers latch superconducting, and the detectors have to be warmed en masse in order to recover them, a process that takes many hours as the focal planes cool again to their operating temperature. Concerned about this possibility, Kyle insisted that they hook the bolometer system up to a set of batteries in order to prevent this in case the power did go out, as is its wont. After frantically assembling a set of cables and going through the switchover procedure, they got the system on batteries -- and not 2 minutes later, the high bay power went out. But the system stayed powered!

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120722

Saturday, July 21, 2012

Day 55, 7/21/2012 -- High and low

In order to manage gondola time more effectively, Franky and I are shifting to an 'early' schedule where we get in to the high bay around 7AM (rather than 9AM as we had been previously). This allows Franky more time to mess with detectors after the end of the night's scanning at ~6AM, and means I get to go up on the tower and bring down our fragile and expensive millimeter-wave source before the temperature starts rising.

So after an invigorating climb up the water tower, I set about doing some gondola work -- I managed to install the set of brackets for the second power box and the battery tables underneath the gondola, necessitating a lot of crawling around on the floor and drilling holes. Once Jeff came in in the afternoon, we also took a set of mirror position measurements to see if they are actually where they should be after having aligned the mirror mounts.

Franky was able to recover a fair number more detector combs, so we have nearly our full complement of detectors available for testing. There are thunderstorms tonight, so there likely won't be any scanning tonight. However, the night crew has some other tests they can do, so it's not a wasted night.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120721

Friday, July 20, 2012

Day 54, 7/20/2012 -- SIPpin' on gin and juice

Today our Science Instrumentation Package, or SIP, was moved into the high bay to prepare for integration with EBEX. The SIP is our interface to the experiment, handing things like command uplink and data downlink (over a variety of different protocols: line-of-sight biphase, TDRSS satellite relay, and Iridium satellite) as well as a small number of high-reliability hardware on/off comands and all of CSBF's payload telemetry. As this is a new SIP, there will be a bit of testing first before it gets installed on the gondola.

After Franky finished investigating some SQUID issues in the morning, we cycled the fridges again (yesterday's cycle didn't last long due to the need to keep the temperature warm for a long period of time for debugging). I made a cable to connect a set of ground batteries to the EBEX bolo power system so we can run the detectors off of cleaner power than the switching power supply that we currently use. I also did a little bit of work on the power system filter box, but found I was missing some connectors Linkthat I need in order to put it together for reals (since ordered!). Britt worked on more parts for the flight power system as well.

The plan for the evening is to take a first cut at doing beam maps. The source will be mounted up on the tower tonight, and the night crew will try and find a detector so they can set levels (to avoid saturation) and hopefully do some first beam maps.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120720

Thursday, July 19, 2012

Day 53, 7/19/2012 -- There! Are! SIX! Lights!

Today we ran the first fridge cycle with the cryostat on the gondola. Things mostly went well (it's pretty awesome to send a single command and have the gondola tilt to the right position and then run the cycle completely automatically) except for a few issues biasing the detectors towards the end. Franky spent the afternoon debugging what went wrong with the detector biasing.

The other big ticket item today was setting up the so-called 'lighthouse' on the water tower, requiring the first of what will be many (many) trips up to the top of the tower. Jeff had the clever idea of using small power bricks to power the lights this year, so instead of lugging two ~25 lb power supplies up the 150 meters of the water tower, we only had to take six ~1 lb (or less) 12VDC power bricks up (along with the rest of the light hardware). KyleH and I got the whole thing set up in about 2 hours, so now we have six lights up on the water tower to use as a 'fake sky' for pointing while doing water tower scans.

On the agenda for tonight is a relative timing test between the ACS and bolo systems as well as scans of the lighthouse to build a source catalog to use in subsequent scans.

Picture (just one -- I'll take more pictures from atop the tower when I don't have to lug other equipment up too):
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120719

Wednesday, July 18, 2012

Day 52, 7/18/2012 -- Water tower scans soon

We are putting the final pieces in place before we are able to start taking calibration measurements with the integrated system.

Jeff and I took a look at the millimeter wave source we use to do a variety of calibration measurements; we were able to couple it to a function generator in order to do time detector constant tests over the full range of frequencies we want to measure. After the time constant tests, the source will then be placed atop the CSBF water tower and aimed down at the high bay and we'll start doing beam maps and our absolute polarization angle calibration. To that end, I also worked out the coordination with Verizon Wireless that they would turn down their cell phone transmitter at our request for the next 3 weeks in the late night/early morning so we can take data without seeing the interference from the tower.

I also managed to route the final dryer hose (containing the cryostat housekeeping signals) into position, and Kyle finished the securing and cable connections. Kyle and Britt also worked all day on checking the system grounding and ensuring the bolo and ACS systems weren't inadvertently shorted together. Britt also did a lot of work on the power system to get it ready to run off batteries when we need to.

Franky spent the day working with the SQUIDs -- the first detector system work that's been done since moving the cryostat to the gondola.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120718

Tuesday, July 17, 2012

Day 51, 7/17/2012 -- Mirror, mirror, on the...gondola?

Today was another insanely productive day. Jeff and I aligned the primary mirror mount in the morning, Franky and others were able to mount all of the detector readout cabling in their metallic 'dryer hoses' (which currently look pretty terrible since we didn't see the point in mangling fresh ones and having to replace them again in Antarctica), Jeff, Michele, Kyle, KyleH, Britt, and I mounted both mirrors, Kyle routed and secured the dryer hoses so they wouldn't interfere when the gondola tilts down in elevation, and Jeff and I installed and measured the position of an inclinometer atop the cryostat to be used in the polarization rotation calibration test. And then Amber's husband grilled dinner for everyone!

Not bad for a day's work.

Cynthia left Palestine today to return to MN to teach a summer course, so we're a little shorthanded at the moment.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120717

Monday, July 16, 2012

Day 50, 7/16/2012 -- Alignment, test fitting, network debugging

Today we had progress on multiple fronts simultaneously. Jeff and I spent a fair bit of today on the gondola atop the cryostat making measurements and adjustments for our mirror mounts. We got the secondary mirror mount aligned and will tackle the primary mirror tomorrow.

Elsewhere in the collaboration, Andrei and KyleH continued to set up hardware for the Bemco test of the liquid cooling system. Kyle plugged in the final set of BRO cables (the gondola network optical fibers), and he and Seth spent the rest of the day debugging network communications issues (successfully).

And last, the Baffle/Triangle Support (BTS) was test-fitted on the gondola along with the triangle spreader bar we use and we verified that everything fits properly (only minor 'convincing' was required to get all of the holes to line up).

I've posted pictures from yesterday in yesterday's post if you want to go back and look at them.

Pictures from today:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120716

Sunday, July 15, 2012

Day 49, 7/15/2012 -- Crises, crates, and cables

I realize that I forgot to mention our little liquid helium crisis back on Friday, so I'll mention it here, because the crisis has been summarily averted:

On Friday, we needed to fill the liquid helium tank of the EBEX cryostat. This was no problem, as we had received a full 500-liter dewar about a week and a half prior that would be sufficient to fill EBEX nearly four times over. The only problem is...it wasn't full. In fact, it wasn't anywhere near full. Actually, it turns out it was completely empty. And warm. We found this out Friday morning, and the helium in EBEX would run out on Sunday if we didn't get our fix. And if the helium ran out, the cryostat would start to warm up...and that's not good at all.

So...we let CSBF know, and they set the wheels in motion to see if we could get helium over the weekend. From their liquid helium vendor...in east Texas. They ended up having two CSBF guys drive the 11-12 hours out to the vendor where a full 500L dewar would await them at noon on Saturday, and then drive all the way back so we would have our helium by midnight Saturday. And...so they did! We were able to fill our helium before the situation got really dire -- CSBF really goes out of their way to make sure we're able to get science done.

Aside from helium, today was largely a continuation of yesterday's work. We started adding the cables for the bolo system to the gondola in order that they could be routed in a usable way without risk of them tearing or being sheared by gondola motion. While doing this, I noticed one of the bolo system power cables was noticeably lighter than the others -- it turns out it had been made with the wrong gauge wire (too small), so I set Cynthia to work on making a new cable while Jeff and I worked out the issues routing the other cables. Jeff also mounted the HWP control and readout crate. Kyle spent most of the day debugging commanding issues in Power Crate 1 (with remote help from Ilan over the phone) and seems to have worked out all of the known problems. So with that done, all of the major components of the bolo system were in place on the gondola with the exception of the readout cables.

Elsewhere in the high bay, Andrei (oh, forgot to mention, he was gone and came back on Thursday) and KyleH started setting up for a Bemco test of our liquid cooling system components. In other comings and goings, Franky arrived from McGill late last night to replace Kevin, who will be leaving tomorrow morning. And Kate and daughter left around mid-day today.

I have pictures, but I forgot to grab the memory card from my camera so their posting will have to wait until tomorrow.

UPDATE -- Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120715

Saturday, July 14, 2012

Day 48, 7/14/2012 -- I gotta say, today was a good day

Some days, everything just goes right. We came in early, broke down the detector readout system, installed the double window parts, removed the cryo mass dummy from the gondola, and installed the cryostat into the gondola -- all in 3 hours. The rest of the day was devoted to mounting the detector readout boxes (the "BROs"), of which we got all 4 mounted, mounting one of the power crates (the other needs some investigations), and tying up some of the cables so Joy and Chappy can do scan tests tonight without having to worry about snagging cables.

All in all, pretty awesome.

And turns out we're not the only ones who have a cryostat mounted on their gondola -- the BLAST team mounted theirs yesterday.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120714